OpenStudy (anonymous):

What is the total differential of w = z*e^(x+y) at (0, 0, 1) ?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

i've seen this before dw=(∂w/∂z)dz+(∂w/∂x)dx+(∂w/∂y)dy

OpenStudy (anonymous):

\[(\partial w/ \partial z)_{(0,0,1)} = e^{0+0} = 1, (\partial w/ \partial x)_{(0,0,1)} = 1. e^{0+0} = 1, (\partial w/\partial y)_{(0,0,1)} = 1. e ^{0+0} = 1\]